# Open-Vocabulary Camouflaged Object Segmentation

### v2

@@ Abstract
Recently, the emergence of the large-scale vision-language model (VLM), such as CLIP, has opened the way towards open-world object perception. Many works have explored the utilization of pre-trained VLM for the challenging open-vocabulary dense prediction task that requires perceiving diverse objects with novel classes at inference time. Existing methods construct experiments based on the public datasets of related tasks, which are not tailored for open vocabulary and rarely involve imperceptible objects camouflaged in complex scenes due to data collection bias and annotation costs. To fill in the gaps, we introduce a new task, open-vocabulary camouflaged object segmentation (OVCOS), and construct a large-scale complex scene dataset (OVCamo) containing 11,483 hand-selected images with fine annotations and corresponding object classes. Further, we build a strong single-stage open-vocabulary camouflaged object segmentation transformer baseline OVCoser attached to the parameter-fixed CLIP with iterative semantic guidance and structure enhancement. By integrating the guidance of class semantic knowledge and the supplement of visual structure cues from the edge and depth information, the proposed method can efficiently capture camouflaged objects. Moreover, this effective framework also surpasses previous state-of-the-arts of open-vocabulary semantic image segmentation by a large margin on our OVCamo dataset. With the proposed dataset and baseline, we hope that this new task with more practical value can further expand the research on open-vocabulary dense prediction tasks. The code and data will be available in the future.
@@
The ability to identify and reason about object regions in a visual scene is essential for a wide range of human activities. As a complex and fundamental task in computer vision, detecting and segmenting objects with diverse appearances in complex scenes is also an important challenge, which is crucial for applications across vision and robotics fields, including autonomous driving [3], medical image analysis [28], and intelligent robotics [62], to name a few. In the past few years, numerous typical methods [19, 7, 6, 54] have emerged with the help of a mass of labeled data, which have greatly promoted the development of related fields such as semantic image segmentation (SIS) [11, 44]. However, existing works have mainly focused on predefined closed-set scenarios, where all semantic concepts are seen during both the inference and training phases. Such a scenario setting oversimplifies the real-world complexity. To this end, many explorations have been contributed to open vocabulary settings [61].
Recently, large-scale pre-trained visual language models (VLMs) such as CLIP [37, 9] have been gaining attention. Image-text matching-based learning mechanism provides them with the ability to align textual and visual signals well, and many works demonstrate their potential for open vocabulary tasks. Besides, data plays an important role in open-vocabulary tasks. Existing open-vocabulary semantic image segmentation (OVSIS) tasks rely on related public datasets [4, 57, 34, 35, 12], while they are not designed for the open-vocabulary setting, and there is high semantic similarity between their class definitions as revealed in [48]. Due to the data collection bias and annotation cost constraints, the existing open-vocabulary benchmarks lack special attention to finely perceive the objects of interest in concealed scenes. And the widely used VLMs are pre-trained on image-text pairs with inherent object concept bias, thus, their ability to segment objects in complex scenes remains to be verified.
@@
In this paper, we introduce a new open-vocabulary segmentation task OVCOS dedicated to analyzing camouflaged object perception in diverse natural scenes. And a large-scale data benchmark, named OVCamo, is carefully constructed. Besides, we also design a strong baseline OVCoser for the proposed OVCOS, based on the VLM-driven single-stage paradigm. The camouflage111To ensure the reliability of the definition, we follow existing published literature and collect data from the camouflage perception field, i.e., CSU [15]. arises from several sources, including similar patterns to the environment (e.g., color and texture) and imperceptible attributes (e.g., small size and heavy occlusion) as statistically illustrated in Fig. 3. Considering the imperceptible appearance of camouflaged objects, accurate recognition and capture actually depend more on the cooperation of multi-source knowledge. As shown in Fig. 1, in addition to visual appearance cues, we introduce the depth for the spatial structure of the scene, the edge for the regional changes about objects, and the text for the context-aware class semantics. Considering the cooperative relationship between class recognition and object perception, the iterative learning strategy is introduced to feed back the optimized semantic relationship, resulting in more accurate object semantic guidance. This top-down conceptual reinforcement can further optimize open-vocabulary segmentation performance. With the help of the iterative multi-source information joint learning strategy, our method OVCoser shows good performance in the proposed OVCOS task.
@@ 2 Related Works
Open Vocabulary Semantic Image Segmentation (OVSIS). Although a variety of pipelines have emerged as summarized in [61], from an overall perspective, their efforts are similar, namely, how to align class name/description semantic embedding with visual features to anchor relevant object cues in the representation space. The early pioneer work [55] attempts to connect word concepts and semantic relations, and encodes word concept hierarchy to parse images. Due to the leading performance of the VLM on the image-text joint modeling, it has been gradually applied in the OVSIS field. In terms of structure, existing schemes can be roughly categorized into two types: two-stage [53, 49, 27, 47] and single-stage [58, 10, 52]. In [53], a rough segmentation map for each class is created from a VLM and then refined by the test-time augmentation. These rough maps are utilized as pseudo-labels for subsequent fine segmentation by stochastic pixel sampling. SimSeg [49] adopts a cascaded design including class-agnostic proposal generation by MaskFormer [7] and class assignment by CLIP [37]. Furthermore, OVSeg [27] fine-tunes CLIP on the noisy but diverse data to improve its generalization to masked images. In lieu of using existing SIS models, a text-to-image diffusion model is introduced to generate mask features with implicit image captions in [47].
The single-stage design is more flexible and simpler. MaskCLIP [58] directly modifies CLIP for semantic segmentation without training, while SAN [48] achieves better performance with the help of adapters. CAT-Seg [10] highlights the importance of the cost aggregation between image and text embeddings for the OVSIS decoding. Recent FC-CLIP [52] investigates the hierarchical CLIP image encoder. Although these methods show impressive OVSIS performance, they still follow the generalized segmentation paradigm [7, 6] in SIS, ignoring valuable auxiliary cues for object perception. This also causes them to struggle with objects camouflaged in complex scenes. Unlike them, our method, which is tailored for OVCOS, can effectively tap into the camouflaged object by integrating task-specific multi-source knowledge from visual appearance, spatial structure, object contour, and class semantics.
@@
Camouflaged Scene Understanding (CSU). This is a research hotspot in the computer vision community, aiming to perceive objects with camouflage [15]. Different from traditional object detection, CSU is obviously a more challenging problem. It can be applied in some specific fields, such as medical analysis [13, 16] and agricultural management [39, 29]. This topic is currently defined as the class-agnostic form, focusing on the area of camouflaged objects in the visual scene. The available work [14, 24, 21, 36, 13, 8, 51] to date has demonstrated promising performance on existing data benchmarks [41, 23, 14, 32, 2, 8]. Unlike previous settings, the proposed OVCOS task requires further perception of object classes. Admittedly, the publicly available data provides critical support for this new task. It helps us take the first step.
@@ 3 OVCamo Dataset
This work focuses on open-vocabulary segmentation in the camouflaged scene. It enriches the connotation of OVSIS and provides a more challenging benchmark.
Image Collection. Our data is collected from existing CSU datasets that have finely annotated segmentation maps. Specifically, the OVCamo integrates 11,483 hand-selected images covering 75 object classes reconstructed from several public datasets [14, 56, 50, 2, 8]. The distribution of the number of samples in different classes is shown in Fig. 2. Meanwhile, we consider attributes of objects when selecting images, such as object concentration, average color ratio, object-image area ratio, number of object parts, and normalized centroid. Tab. 1 gives their definitions. And Fig. 3 visualizes the attribute distribution of the proposed dataset. The camouflaged objects of interest usually have complex shape Fig. 2(a), high similarity to the background Fig. 2(b), and small size Fig. 2(c). And the image often contains multiple camouflaged objects or sub-regions with a central bias as shown in Fig. 2(d) and Fig. 2(e).
| Attribute |
Description |
| Object Concentration |
Object pixel concentration, which calculates the area ratio between the object region and its minimum rotatable bounding box. |
| Average Color Ratio |
Color ratio of the object to the background, which calculates the average of the three color channels in their respective regions. |
| Object-Image Area Ratio |
Area ratio of the object relative to the image. |
| Number of Object Parts |
Number of separate areas in the image that belong to the object. |
| Normalized Centroid |
Object centroid coordinates, which are normalized using the image shape. |
Re-annotation. The original annotation cannot be used directly in the open-vocabulary setting, due to the following semantic ambiguities caused by different annotation standards, i.e., 1) Broad concepts, such as “fish” and “bug”; 2) Vague definitions, such as “small fish” and “black cat”; 3) Inconsistent granularity, such as the coexistence of “orchid mantis” and “mantis”; 4) Non-entity concepts, such as “other”. These issues can lead to unreasonable and unreliable results for the open-vocabulary prediction as discussed in [59]. To this end, we relabel the classes of all camouflaged objects and take the generality of the concept as the criterion for class definition, which also ensures lower semantic similarity.
@@ 4.1 Overall Architecture
We follow the common encoder-decoder paradigm and the pipeline is shown in Fig. 4. Specifically, we first leverage the textual encoder of the frozen CLIP to extract semantic embedding from the class label set , and the visual encoder to extract multi-scale image features . Both of them are fed into the decoder as the information bedrock of object segmentation, as shown in Fig. 5. Structural cues such as depth and edge are also introduced to assist the iterative refinement process. Finally, the class-related segmentation which is the segmentation logits after the sigmoid function processing, is used to remove the interference from the background in the high-level image feature and guide the generation of object-oriented visual representation . And the class label is determined by the similarity matching between and .
Details of and . In the proposed model, the visual encoder and embedding layer together are used to extract the high-level embedding corresponding to the object of interest in the input image. The two independent sub-networks are split from the visual network of CLIP [37, 9]. contains all the feature encoding layers for extracting the multi-scale image features. And corresponds to the final high-dimensional projection layer, which is used to convert the high-level image feature into the visual embedding vector .
4.2 Semantic Guidance (SG)
The class label definition usually is independent of the image scene. It is very important to fully utilize the class prior for object recognition in complex scenes. In the proposed decoder, normalized textual embedding is introduced into each stage to highlight semantically relevant cues. As shown in Fig. 5(a), we design a semantic guidance component SG to inject concept cues into the self-enhancement of the image feature.
@@ 4.2 Semantic Guidance (SG)
Semantic Guidance Attention (SGA). Specifically, the normalized image feature is linearly mapped to , , and , while the textual embedding is transformed to the class guidance vector . The similarity between and reflects the activation of different classes in spatial locations. In Agg of Fig. 5(a), the base weight for the spatial guidance is obtained after highlighting the most relevant class information by softmax operation. And then is modulated and fed into MHSA, i.e., multi-head self-attention [45].
4.3 Structure Enhancement (SE)
@@
Existing methods demonstrate that low-level structural information, such as the edge [43] and the depth [46], plays an important role in CSU, which is closely related to the mechanism of the human visual system. So the SE component attached to the low-level SG is proposed to integrate the edge-aware and depth-aware cues and improve the structural details. Specifically, the output of the SG is fed into two separate branches containing the convolutional stem and head for the edge and depth estimation as shown in Fig. 5. The edge and depth logits maps, i.e., and , from the head in the branch of the layer are directly supervised. And the outputs and of the stem are fed into the SE. In the structure enhancement attention (SEA), they independently update the normalized visual features using MHSA [45], and the corresponding outputs are combined with the learnable weight as in Fig. 5(b).
4.4 Iterative Refinement
@@
In the SG component, the aggregation process between image features and class semantics is not aligned and requires data-driven optimization. Considering the aligned embedding space of the pre-trained CLIP, we introduce the correlation matrix between the visual and textual embeddings into the SG as shown in Fig. 5(a). Meanwhile, due to the emphasis of the decoder output on the object region, the object-aware representation is also inputted, which comes from the image features pooled by the coarse segmentation prediction in the last iteration. By combining the two, we obtain task-oriented object cues, which is actually inspired by the top-down attention mechanism in the human cognitive system [22, 1]. The spatial activation map of such object cues over image features is used to re-modulate . Besides, the SE in the iteration also helps the model to further optimize the texture details. To benefit as much as possible from the assistance from structure enhancement while avoiding over-computation, we set the iteration entry to the third decoding layer as shown in Fig. 5.
4.5 CamoPrompts
@@
As mentioned in [37], prompt engineering and ensembling are important for the transfer performance of CLIP on downstream tasks, and the prompt template should be more relevant to the data type. Because additional task-related cues are generally able to impose the necessary contextual constraint to the flexible CLIP. Hence, instead of common practices [60, 37, 18], we design a simpler yet more effective template set CamoPrompts tailored for OVCOS to decorate the class name, and average their textual embeddings as the final semantic embedding for each class. Its full form is depicted in Fig. 8, while it also achieves better classification performance in comparison with other forms.
4.6 Supervision
@@
In each iteration, in addition to semantic object segmentation, we also need to perform depth estimation and edge estimation as auxiliary tasks. For the segmentation prediction, we follow the commonly used weighted segmentation loss function [51, 43]. For the edge estimation, considering the imbalance problem of positive and negative samples, we introduce the dice loss function as . The summation of L1 and SSIM losses, i.e., , is used for the depth estimation. The total loss of our method can be formulated as follows:
where and are used to index iterations and layers, respectively. And the total number of iterations is set to 2 as mentioned in Sec. 5.3.

### v3

@@ Abstract
Recently, the emergence of the large-scale vision-language model (VLM), such as CLIP, has opened the way towards open-world object perception. Many works have explored the utilization of pre-trained VLM for the challenging open-vocabulary dense prediction task that requires perceiving diverse objects with novel classes at inference time. Existing methods construct experiments based on the public datasets of related tasks, which are not tailored for open vocabulary and rarely involve imperceptible objects camouflaged in complex scenes due to data collection bias and annotation costs. To fill in the gaps, we introduce a new task, open-vocabulary camouflaged object segmentation (OVCOS), and construct a large-scale complex scene dataset (OVCamo) containing 11,483 hand-selected images with fine annotations and corresponding object classes. Further, we build a strong single-stage open-vocabulary camouflaged object segmentation transformer baseline OVCoser attached to the parameter-fixed CLIP with iterative semantic guidance and structure enhancement. By integrating the guidance of class semantic knowledge and the supplement of visual structure cues from the edge and depth information, the proposed method can efficiently capture camouflaged objects. Moreover, this effective framework also surpasses previous state-of-the-arts of open-vocabulary semantic image segmentation by a large margin on our OVCamo dataset. With the proposed dataset and baseline, we hope that this new task with more practical value can further expand the research on open-vocabulary dense prediction tasks. Our code and data can be found in the link.
@@
The ability to identify and reason about object regions in a visual scene is essential for a wide range of human activities. As a complex and fundamental task in computer vision, detecting and segmenting objects with diverse appearances in complex scenes is also an important challenge, which is crucial for applications across vision and robotics fields, including autonomous driving [3], medical image analysis [35], and intelligent robotics [90], to name a few. In the past few years, numerous typical methods [19, 7, 6, 75] have emerged with the help of a mass of labeled data, which have greatly promoted the development of related fields such as semantic image segmentation (SIS) [11, 57]. However, existing works have mainly focused on predefined closed-set scenarios, where all semantic concepts are seen during both the inference and training phases. Such a scenario setting oversimplifies the real-world complexity. To this end, many explorations have been contributed to open vocabulary settings [89, 60]. In existing work, open vocabulary learning utilizes visual-related language data (e.g., object class text or descriptions) to align images and language features, enabling models to perceive novel classes without extensive labeled data [60]. Recently, large-scale pre-trained visual language models (VLMs) such as CLIP [50, 9] have been gaining attention. Image-text matching-based learning mechanism provides them with the ability to align textual and visual signals well, and many works demonstrate their potential for open vocabulary tasks. Besides, data plays an important role in open-vocabulary tasks. Existing open-vocabulary semantic image segmentation (OVSIS) tasks rely on related public datasets [4, 85, 43, 44, 12], while they are not designed for the open-vocabulary setting, and there is high semantic similarity between their class definitions as revealed in [63]. Due to the data collection bias and annotation cost constraints, the existing open-vocabulary benchmarks lack special attention to finely perceive the objects of interest in concealed scenes. And the widely used VLMs are pre-trained on image-text pairs with inherent object concept bias, thus, their ability to segment objects in complex scenes remains to be verified.
@@
In this paper, we introduce a new open-vocabulary segmentation task OVCOS dedicated to analyzing camouflaged object perception in diverse natural scenes. And a large-scale data benchmark, named OVCamo, is carefully constructed. Besides, we also design a strong baseline OVCoser for the proposed OVCOS, based on the VLM-driven single-stage paradigm. The camouflage111To ensure the reliability of the definition, we follow existing published literature and collect data from the class-agnostic camouflage perception field, i.e., CSU [15]. arises from several sources, including similar patterns to the environment (e.g., color and texture) and imperceptible attributes (e.g., small size and heavy occlusion) as statistically illustrated in Fig. 4. Considering the imperceptible appearance of camouflaged objects, accurate recognition and capture actually depend more on the cooperation of multi-source knowledge. As shown in Fig. 1, in addition to visual appearance cues, we introduce the depth for the spatial structure of the scene, the edge for the regional changes about objects, and the text for the context-aware class semantics. Considering the cooperative relationship between class recognition and object perception, the iterative learning strategy is introduced to feed back the optimized semantic relationship, resulting in more accurate object semantic guidance. This top-down conceptual reinforcement can further optimize open-vocabulary segmentation performance. With the help of the iterative multi-source information joint learning strategy, our method OVCoser shows good performance in the proposed OVCOS task.
@@ 2 Related Works
Open Vocabulary Semantic Image Segmentation (OVSIS). Although a variety of pipelines have emerged as summarized in [89], from an overall perspective, their efforts are similar, namely, how to align class name/description semantic embedding with visual features to anchor relevant object cues in the representation space. The early pioneer work [76] attempts to connect word concepts and semantic relations, and encodes word concept hierarchy to parse images. Due to the leading performance of the VLM on the image-text joint modeling, it has been gradually applied in the OVSIS field. In terms of structure, existing schemes can be roughly categorized into two types: two-stage [71, 64, 33, 62] and single-stage [86, 10, 70]. In [71], a rough segmentation map for each class is created from a VLM and then refined by the test-time augmentation. These rough maps are utilized as pseudo-labels for subsequent fine segmentation by stochastic pixel sampling. SimSeg [64] adopts a cascaded design including class-agnostic proposal generation by MaskFormer [7] and class assignment by CLIP [50]. Furthermore, OVSeg [33] fine-tunes CLIP on the noisy but diverse data to improve its generalization to masked images. In lieu of using existing SIS models, a text-to-image diffusion model is introduced to generate mask features with implicit image captions in [62]. The single-stage design is more flexible and simpler. MaskCLIP [86] directly modifies CLIP for semantic segmentation without training, while SAN [63] achieves better performance with the help of adapters. CAT-Seg [10] highlights the importance of the cost aggregation between image and text embeddings for the OVSIS decoding. Recent FC-CLIP [70] investigates the hierarchical CLIP image encoder. Although these methods show impressive OVSIS performance, they still follow the generalized segmentation paradigm [7, 6] in SIS, ignoring valuable auxiliary cues for object perception. This also causes them to struggle with objects camouflaged in complex scenes. Unlike them, our method, which is tailored for OVCOS, can effectively tap into the camouflaged object by integrating task-specific multi-source knowledge from visual appearance, spatial structure, object contour, and class semantics.
@@
Camouflaged Scene Understanding (CSU). This is a research hotspot in the computer vision community, aiming to perceive objects with camouflage [15]. Different from traditional object perception like salient object detection [22, 73, 30, 72, 23, 74, 31, 48, 45, 83, 81, 80, 69, 78, 79, 49, 82] and semantic object segmentation [21, 77], CSU is obviously a more challenging problem. It can be applied in some specific fields, such as medical analysis [13, 16] and agricultural management [52, 36]. This topic is currently defined as the class-agnostic form, focusing on the area of camouflaged objects in the visual scene. The available work [14, 28, 24, 46, 47, 13, 8, 68] to date has demonstrated promising performance on existing data benchmarks [54, 27, 14, 40, 2, 8]. Unlike previous settings, the proposed OVCOS task requires further perception of object classes. Admittedly, the publicly available data provides critical support for this new task. It helps us take the first step.
@@ 3 OVCamo Dataset
This work focuses on open-vocabulary segmentation in the camouflaged scene. It enriches the connotation of OVSIS and provides a more challenging benchmark. Note that our aim is not to claim novelty through class annotations alone, but rather to focus on how such annotations can pave the way for new tasks and methods. Our dataset provides additional features such as clean data, less ambiguity, and sensible class hierarchies, which provide the necessary data for solving the new and complex task. It not only adds complexity and challenge, but also promotes innovation and advances the field through the need for advanced methods. We believe that this dataset can lead to new opportunities in areas such as wildlife conservation and intelligence sensing.
| Attribute | Description |
| Object Concentration | Object pixel concentration, which calculates the area ratio between the object region and its minimum rotatable bounding box. |
| Average Color Ratio | Color ratio of the object to the background, which calculates the average of the three color channels in their respective regions. |
| Object-Image Area Ratio | Area ratio of the object relative to the image. |
| Number of Object Parts | Number of separate areas in the image that belong to the object. |
| Normalized Centroid | Object centroid coordinates, which are normalized using the image shape. |
Image Collection. Our data is collected from existing CSU datasets that have finely annotated segmentation maps. Specifically, the OVCamo integrates 11,483 hand-selected images covering 75 object classes reconstructed from several public datasets [14, 84, 65, 2, 8]. The distribution of the number of samples in different classes is shown in Fig. 2. Meanwhile, we consider attributes of objects when selecting images, such as object concentration, average color ratio, object-image area ratio, number of object parts, and normalized centroid. Tab. 1 gives their definitions. And Fig. 4 visualizes the attribute distribution of the proposed dataset. The camouflaged objects of interest usually have complex shape Fig. 4(a), high similarity to the background Fig. 4(b), and small size Fig. 4(c). And the image often contains multiple camouflaged objects or sub-regions with a central bias as shown in Figs. 4(d) and 4(e).
Annotation. The original annotation cannot be used directly in the open-vocabulary setting, due to the following semantic ambiguities caused by different annotation standards, i.e., 1) Broad concepts, such as “fish” and “bug”; 2) Vague definitions, such as “small fish” and “black cat”; 3) Inconsistent granularity, such as the coexistence of “orchid mantis” and “mantis”; 4) Non-entity concepts, such as “other”. These issues can lead to unreasonable and unreliable results for the open-vocabulary prediction as discussed in [87]. To this end, we relabel the classes of all camouflaged objects and take the generality of the concept as the criterion for class definition, which also ensures lower semantic similarity. The similarity analysis and rough hierarchy of the class set can be found in Appendix 0.B.
@@ 4 Methodology
Overall Architecture. We follow the common encoder-decoder paradigm and the pipeline is shown in Fig. 5. Specifically, we first leverage the textual encoder of the frozen CLIP to extract semantic embedding from the class label set , and the visual encoder to extract multi-scale image features . Both of them are fed into the decoder as the information bedrock of object segmentation, as shown in Fig. 6. Structural cues such as depth and edge are also introduced to assist the iterative refinement process. Finally, the class-related segmentation which is the segmentation logits after the sigmoid function processing, is used to remove the interference from the background in the high-level image feature and guide the generation of object-oriented visual representation . And the class label is determined by the similarity matching between and . More details about the proposed architecture are listed in Appendix 0.A.
@@
Semantic Guidance (SG). The class label definition usually is independent of the image scene. It is very important to fully utilize the class prior for object recognition in complex scenes. In the proposed decoder, normalized textual embedding is introduced into each stage to highlight semantically relevant cues. As shown in Fig. 7(a), we design a semantic guidance component SG to inject concept cues into the self-enhancement of the image feature. Specifically, in our semantic guidance attention (SGA), the normalized image feature is linearly mapped to , , and , while the textual embedding is transformed to the class guidance vector . The similarity between and reflects the activation of different classes in spatial locations. In Agg of Fig. 7(a), the base weight for the spatial guidance is obtained after highlighting the most relevant class information by softmax operation. And then is modulated and fed into MHSA, i.e., multi-head self-attention [59].
@@
Structure Enhancement (SE). Existing methods demonstrate that low-level structural information, such as the edge [56] and the depth [61], plays an important role in CSU, which is closely related to the mechanism of the human visual system. So the SE component attached to the low-level SG is proposed to integrate the edge-aware and depth-aware cues and improve the structural details. Specifically, the output of the SG is fed into two separate branches containing the convolutional stem and head for the edge and depth estimation as shown in Fig. 6. The edge and depth logits maps, i.e., and , from the head in the branch of the layer are directly supervised. And the outputs and of the stem are fed into the SE. In the structure enhancement attention (SEA), they independently update the normalized visual features using MHSA [59], and the corresponding outputs are combined with the learnable weight as in Fig. 7(b).
@@
Iterative Refinement. In the SG component, the aggregation process between image features and class semantics is not aligned and requires data-driven optimization. Considering the aligned embedding space of the pre-trained CLIP, we introduce the correlation matrix between the visual and textual embeddings into the SG as shown in Fig. 7(a). Meanwhile, due to the emphasis of the decoder output on the object region, the object-aware representation is also inputted, which comes from the image features pooled by the coarse segmentation prediction in the last iteration. By combining the two, we obtain task-oriented object cues, which is actually inspired by the top-down attention mechanism in the human cognitive system [25, 1]. The spatial activation map of such object cues over image features is used to re-modulate . Besides, the SE in the iteration also helps the model to further optimize the texture details. To benefit as much as possible from the assistance from structure enhancement while avoiding over-computation, we set the iteration entry to the third decoding layer as shown in Fig. 6.
@@
CamoPrompts. As mentioned in [50], prompt engineering and ensembling are important for the transfer performance of CLIP on downstream tasks, and the prompt template should be more relevant to the data type. Because additional task-related cues are generally able to impose the necessary contextual constraint to the flexible CLIP. Hence, instead of common practices [88, 50, 18], we design a simpler yet more effective template set CamoPrompts tailored for OVCOS to decorate the class name, and average their textual embeddings as the final semantic embedding for each class. Its full form is depicted in Sec. 5.3, while it also achieves better classification performance in comparison with other forms.
@@
Supervision. In each iteration, in addition to semantic object segmentation, we also need to perform depth estimation and edge estimation as auxiliary tasks. For the segmentation prediction, we follow the commonly used weighted segmentation loss function [68, 56]. For the edge estimation, considering the imbalance problem of positive and negative samples, we introduce the dice loss function as . The summation of L1 and SSIM losses, i.e., , is used for the depth estimation. The total loss of our method can be formulated as follows: , where and are used to index iterations and layers, respectively. And the total number of iterations is set to 2 as mentioned in Sec. 5.3.